OpenStudy (masterchief):

what is the inverse function of f(x)= (x + 2)3

OpenStudy (moldybubblegum11):

This is how you find Inverses Step 1 y = x3 + 2. Step 2 x = y3 + 2. Step 3 x - 2 = y3. (x - 2)^(1/3) = y. Step 4 f-1(x) = (x - 2)^(1/3).

OpenStudy (moldybubblegum11):

sorry it took so long